OpenStudy (anonymous):

What is the equation of a line passing through (-3, 7) and having a slope of -1/5?

pooja195 (pooja195):

In point slope form?

OpenStudy (anonymous):

yes

pooja195 (pooja195):

\[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1)

OpenStudy (anonymous):

right....

pooja195 (pooja195):

Can you plug in the given info? :) And let me know what you get.

OpenStudy (anonymous):

y-7=-1/5(x--3)

pooja195 (pooja195):

y-7=-1/5(x--3) - + - = + \[\huge~\rm~y-7=-1/5(x\bf{+}3) \]

OpenStudy (anonymous):

ok, so now what d i do?

pooja195 (pooja195):

Thats all =) that is point slope form